Please show step by step

Convert from degrees Kelvin to degrees Celsius

Formula: ° C = ° K- 273 °
328 ° K = _____ ° C

623 ° K = _____ ° C


Convert from degrees Celsius to degrees Kelvin

Formula: ° K = ° C + 273 °
123 ° C = _____ ° K

39 ° C = _____ ° K


Convert from degrees Celsius to degrees Fahrenheit

Formula: ° F = 9/5 ° C + 32 °
73 ° C = _____ ° F

32 ° C = _____ ° F


Convert from degrees Fahrenheit to degrees Celsius

Formula: ° C = 5/9 (° F- 32 °)
104 ° F = _____ ° C

212 ° F = _____ ° C
Answer #1

°C K - 2730 328 k = 328-273 3.555 ; 623°k Logos = . . 623-273 = 350 °C Ok 3 °C # 273 123°c - 123+ 273 = 396K 39°c = 3a+273 =212 °F » I (2012-32) 5x180 100°c